ROTATIONALLY RESOLVED PHOTOELECTRON SPECTROSCOPY STUDY OF THE FIRST ELEC-TRONIC STATES OF Ar_{2}^{+}

\underline{PATRICK RUPPER} and FRÉDÉRIC MERKT, Physical Chemistry Laboratory, ETH Zurich, CH-8093 Zurich.

The I(1/2u), I(3/2g), I(1/2g), I(3/2u) and II(1/2u) states of Ar_{2}^{+} have been investigated in the region between 124650 cm^{−1} and 128150 cm^{−1} by pulsed-field-ionization zero-kinetic-energy (PFI-ZEKE) photoelectron spectroscopy following (1 + 1′) resonance-enhanced two-photon excitation via the 0_{u}^{+} Rydberg state located below the Ar^{∗} ([3p]^{5}4s′[1/2]_{1}) + Ar (^{1}S_{0}) dissociation limit of Ar_{2}. By selecting single rotational levels of the intermediate state, the rotational structure of four of the six lowest ionic states of Ar_{2}^{+} could be observed and unambiguous assignments of electronic symmetry of the ionic states could be made on the basis of photoionization selection rules. The rotational structure also provided information on the equilibrium internuclear distances R_{e}^{+} for the I(1/2u), I(3/2g), I(1/2g) and II(1/2u) states (R_{e}^{+}(I(1/2u)) = (2.393±0.043) Å, R_{e}^{+}(I(3/2g)) = (2.993 ± 0.005) Å, R_{e}^{+}(I(1/2g)) = (3.151 ± 0.002) Å and R_{e}^{+}(II(1/2u)) = (3.851 ± 0.012) Å). The adiabatic ionization potentials are determined to be IP(I(3/2g)) = (125685.3 ± 0.7) cm^{−1}, IP(I(1/2g)) = (126884.1 ± 0.7) cm^{−1}, IP(I(3/2u)) = (127044.6 ± 1.5) cm^{−1} and IP(II(1/2u)) = (128000.8 ± 0.7) cm^{−1}, from which dissociation energies are obtained as D_{0}^{+}(I(3/2g)) = (1509.3±1.2) cm^{−1}, D_{0}^{+}(I(1/2g)) = (310.5±1.2) cm^{−1}, D_{0}^{+}(I(3/2u)) = (150.0±1.8) cm^{−1} and D_{0}^{+}(II(1/2u)) = (625.4 ± 1.2) cm^{−1}. The interaction between the I(1/2g) and I(3/2g) states could also be observed and analyzed.